OpenStudy (amberlykhan):

Hallie subtracted a quantity from the polynomial 3y^2 + 8y - 16 and produced the expression (y + 2)(y - 2). What quantity did Hallie subtract? Explain how you got your answer.

OpenStudy (calculusxy):

I will try to do the work with you... We have the starting polynomial (the one that we will subtract from) to be \(3y^2 + 8y - 16\) and the unknown to be \(x\). The difference will be \((y+2)(y - 2)\). \(3y^2 + 8y - 16 - x = (y + 2)(y - 2)\). I would suggest for you to expand from the factorized form of \((y+ 2)(y - 2)\) by using \(\text {FOIL}\).

OpenStudy (calculusxy):

Make \((y+2)(y-2)\) a polynomial by using FOIL. Do you know how to do that?

OpenStudy (amberlykhan):

Yes. After doing that I would set it equal to the one you mentioned, and then I would simplify, right?

OpenStudy (calculusxy):

If you mean to isolate \(x\) and find the value of it, then yes.

OpenStudy (amberlykhan):

Oh yeah, I got it. Thanks! I'll let you know what I got just to make sure.

OpenStudy (calculusxy):

Okay

OpenStudy (calculusxy):

@AmberlyKhan Did you get it?

OpenStudy (amberlykhan):

Is it x = 12 + 2y^2 - 8y?

OpenStudy (amberlykhan):

And I would factor that?

OpenStudy (amberlykhan):

I did something wrong, because the difference I got was y^2 - 28. It shoud be y^2 - 4.

OpenStudy (calculusxy):

Let's check: \((y+2)(y-2) = y^2 - 2y + 2y - 4 = y^2 - 4\) \((3y^2 + 8y - 16) - x = y^2 - 4\) \((3y^2 + 8y - 16) - (3y^2 + 8y - 16) - x = (y^2 - 4) - (3y^2 + 8y - 16)\) \(-x = y^2 - 4 - 3y^2 - 8y + 16\) \(-x = - 2y^2 - 8y + 12\) \(x = 2y^2 + 8y - 12\)

OpenStudy (amberlykhan):

How'd you get +16? I got -16.

OpenStudy (calculusxy):

There's an invisible -1 in front of \((3y^2 + 8y - 16)\). So it's basically like you're distributing -1 to -16 to get 16.

OpenStudy (amberlykhan):

Oh, nevermind, I got it.

OpenStudy (amberlykhan):

You distributed the negative.

OpenStudy (calculusxy):

Yes I distributed the -1

OpenStudy (calculusxy):

Do you understand this now?

OpenStudy (amberlykhan):

OK, I got the answer. Thanks for bearing with me!